five years ago akbar was 8 times as old as his son ten years later he will be three times as old as his son what is akbars present age?
Answer by lwsshak3(11628)

five years ago akbar was 8 times as old as his son. ten years later he will be three times as old as his son what is akbars present age?
***
5 yrs ago:
x=son's age
8x=akbar's age
..
10 yrs later:
x+10=son's age
8x+10=akbars age
8x+10=3(x+10)=3x+30
5x=20
x=4
8x=32 (ackbar's age 5 yrs ago)
..
akbar's present age=32+5=37